you are standing in line at ticket window. there are 5 more people behind you then ahead of you.the total number of people
in line including you is three times the number of people ahead of you . How many people are ahead of you

Answer by ikleyn(52908)

.
you are standing in line at ticket window. there are 5 more people behind you then ahead of you.the total number of people
in line including you is three times the number of people ahead of you . How many people are ahead of you
~~~~~~~~~~~~~~~~~~~~~~~~~~~

Let b = no. of people behind you,
    a = no. of people ahead  you.

The condition says that

b - a     = 5       (1)     ("there are 5 more people behind you then ahead of you.")
b + a + 1 = 3a      (2)     ("the total number of people in line including you is three times the number of people ahead of you")


Subtract and solve for "a":
2a+1 = 3a-5
a = 6 

Anser.  There are 6 people ahead of you.